# spacing=nonfrench

### v1

@@ Abstract
The SSH protocol provides secure access to network services, particularly remote terminal login and file transfer within organizational networks and to over 15 million servers on the open internet. SSH uses an authenticated key exchange to establish a secure channel between a client and a server, which protects the confidentiality and integrity of messages sent in either direction. The secure channel prevents message manipulation, replay, insertion, deletion, and reordering. At the network level, SSH uses the SSH Binary Packet Protocol over TCP.
In this paper, we show that as new encryption algorithms and mitigations were added to SSH, the SSH Binary Packet Protocol is no longer a secure channel: SSH channel integrity (INT-PST) is broken for three widely used encryption modes. This allows prefix truncation attacks where some encrypted packets at the beginning of the SSH channel can be deleted without the client or server noticing it. We demonstrate several real-world applications of this attack. We show that we can fully break SSH extension negotiation (RFC 8308), such that an attacker can downgrade the public key algorithms for user authentication or turn off a new countermeasure against keystroke timing attacks introduced in OpenSSH 9.5. We also identified an implementation flaw in AsyncSSH that, together with prefix truncation, allows an attacker to redirect the victim’s login into a shell controlled by the attacker.
In an internet-wide scan for vulnerable encryption modes and support for extension negotiation, we find that 77% of SSH servers support an exploitable encryption mode, while 57% even list it as their preferred choice.
We identify two root causes that enable these attacks: First, the SSH handshake supports optional messages that are not authenticated. Second, SSH does not reset message sequence numbers when encryption is enabled. Based on this analysis, we propose effective and backward-compatible changes to SSH that mitigate our attacks.
@@ 1 Introduction
While TLS is commonly used to secure user-facing protocols such as web, email, or FTP, SSH is used by administrators to deploy and maintain these servers, often with high privilege (root) access and a large attack surface for lateral movement within an organization’s infrastructure. As such, SSH was developed by Tatu Ylonen in 1995 as a secure alternative to telnet and rlogin/rcp and has since become a critical component of internet security.
@@
An SSH connection between a client and a server begins with the Transport Layer Protocol [29], which defines the handshake messages for key exchange and server authentication and how messages are exchanged over TCP using the SSH Binary Packet Protocol (BPP). After the handshake, SSH provides a secure channel for application data. At the application level, the client chooses a sequence of services to run. In practice, the client will run precisely two services: the Authentication Protocol [26] for user authentication with a password or public key, followed by the Connection Protocol [27] for the bulk of SSH’s features like terminal sessions, port forwarding, and file transfer.
@@ 1.1 SSH Channel Security
Figure 1: SSH handshake using a finite-field Diffie-Hellman key exchange. Included sequence numbers are implicit and maintained by the BPP. $\mathsf{Snd}$ denotes the counter for sent packets and $\mathsf{Rcv}$ for received packets. Sequence numbers verified using authenticated encryption are in bold.
In this work, we focus on the integrity of the SSH handshake and the resulting secure channel, as shown in Figure 1. After an initial exchange of version information directly over TCP, the BPP exchanges packets, each containing precisely one message. Initially, the BPP is used without encryption or authentication for the duration of the handshake until the NewKeys message. Afterward, the encryption and authentication keys are used to form a secure channel, with the intent to protect the confidentiality and integrity of the ordered stream of all following messages. Note that technically, the secure channel consists of two separate cipher streams, one for each direction, and that the order of message arrival is only guaranteed for each direction separately.
@@
As SSH is an interactive protocol, the integrity of each packet must be verified when it is received so that it can be promptly decrypted and processed. For this, the BPP appends a Message Authentication Code (MAC) to each packet. A cipher mode and a MAC form an authenticated encryption scheme [4]. SSH historically uses Encrypt-and-MAC (EaM), where the MAC is computed over the plaintext, but this is vulnerable to oracle attacks [2]. Later, Encrypt-then-MAC (EtM) was added, where the MAC is computed over the ciphertext instead. SSH has recently adopted the AEAD ciphers AES-GCM and ChaCha20-Poly1305, where ciphertext integrity is built into the encryption scheme [38].
@@ A Trivial Example: Suffix Truncation Attacks
Note that a per-packet MAC cannot fully protect the channel’s integrity, as packets are verified and decrypted before the end of communication has been seen. This allows for a trivial suffix truncation attack, where the attacker interrupts the message flow at some point during the communication. This is an inherent limitation of interactive protocols and an accepted trade-off in the design of SSH, but also, e.g., the TLS Record Layer. Although this attack cannot be prevented, it can be detected by requiring “end-of-communication” messages as the last messages in both directions. Unfortunately, the SSH standard does not define such messages, while TLS defines “close_notify” alerts for this purpose [37]. For SSH, it is left to the user to notice suspicious interruptions of SSH sessions.
@@
If the MAC was only computed over the payload of each packet, an attacker could still delete, replay, or reorder packets. Therefore, a sequence number is included in the MAC computation, corresponding to the position of the message in the stream. Each peer maintains two distinct counters (starting at 0), one for each direction. The $\mathsf{Snd}$ counter is incremented after a packet has been sent, and the $\mathsf{Rcv}$ counter is incremented after a received packet has been processed. Once the secure channel has been established, the current value of the $\mathsf{Snd}$ counter is used to compute the MAC of an outgoing packet and the current $\mathsf{Rcv}$ to verify the MAC of an incoming packet. If packets in the secure channel are deleted, replayed, or reordered, the sequence numbers get out of sync, and MAC verification will fail.
@@
For TLS, the security guarantees of the Record Layer were formalized as stateful length-hiding encryption [35], where the state mainly consists of the implicit sequence number. The security of the BPP and implicit sequence numbers was analyzed by Bellare et al. in [3] and later refined and extended by Paterson and Watson [36] and Albrecht et al. [1]. These works define, in slightly idealized scenarios, the following informal security goal for a secure channel:
When a secure channel between A and B is used, the data stream received by B should be identical to the one sent by A and vice versa (INT-PST, [15]).
Within their idealization, all three works confirm that the BPP is indeed a secure channel. The difference between the models is that [36] also includes the encrypted length fields of the CBC Encrypt-and-MAC modes, and [1] considers the more recent cipher modes ChaCha20-Poly1305, AES-GCM, and generic Encrypt-then-MAC. Our attacks show that the models underlying the proofs in [1] are only partially accurate. We will explain the discrepancies between the proofs and our findings in Section 2.
@@ 1.2 Overview of Our Attacks on SSH
We consider a MitM attacker who can observe, change, delete or insert bytes at the TCP layer. We assume that the attacker can not break the confidentiality of the session keys, i.e. the attacker has no information about the derived encryption keys, MAC keys, or IV.
@@
Based on these two key observations, we present a series of novel attacks on SSH with increasing complexity and impact.
@@
We show that an attacker can increase the receive counters of the server and the client by inserting messages into the handshake (Section 4.1). Although not required for any of our attacks, we also show that, for some implementations, an attacker can fully control the receive and send counters, setting them to arbitrary, attacker-chosen values (Appendix A).
@@ A Prefix Truncation Attack on the BPP
An attacker can use sequence number manipulation to delete a chosen number of packets at the beginning of the secure channel. Neither the client nor the server detects this change. This breaks the channel integrity of SSH (Section 4.2).
@@
As a practical example, we show an attack that uses prefix truncation to break extension negotiation [9], thereby downgrading the security of the connection. The attacked client might mistakenly believe that the server does not support recent signature algorithms for user authentication or does not implement certain countermeasures to attacks (Section 5.2). Specifically, we can turn off protection against keystroke timing attacks in the recently released OpenSSH 9.5.
@@
As another example, we show two attacks on the AsyncSSH client and server. In the first attack, the victim’s extension info message is replaced with one chosen by the attacker (Section 6.1). For the second attack, the attacker must have a user account on the same server as the victim. The attacker injects a malicious user authentication message so that the victim logs into a shell controlled by the attacker rather than the victim’s shell, thereby giving the attacker complete control over the victim’s terminal input (Section 6.2). These attacks rely on implementation flaws of the AsyncSSH server, in addition to the prefix truncation attack.
@@
Our attacks critically depend on the SSH encryption mode negotiated between the client and the server. The attack works best with the AEAD cipher ChaCha20-Poly1305 (added in 2013). The attack also works with significant probability with any CBC-EtM mode (added in 2012), while CTR-EtM is vulnerable but not exploitable. On the other hand, CBC-EaM, CTR-EaM, and GCM modes are not vulnerable. See Section 4.3 for a complete analysis.
In an internet-wide scan, we show that despite these limitations, 77% of all SSH servers on the internet support a vulnerable encryption mode, and 57% even list it as their preferred choice (Section 7).
@@ 1.3 Our Contributions
  • A first security analysis of SSH extension negotiation, including a novel downgrade attack that disables extension negotiation completely. Consequently, support for public key signature algorithms or, in the case of OpenSSH 9.5, protection against keystroke timing attacks is disabled.
@@
Proof-of-concept implementations for our attacks and the results of our internet-wide scan in an aggregated form are available under the Apache-2.0 open-source license. See:
@@
We started a coordinated disclosure and are in discussion with the maintainers of OpenSSH and AsyncSSH, who acknowledge our findings. Other vendors of major SSH implementations will be contacted later in the process. For our internet-wide scans, we provide an opt-out option and an email address for inquiries. Additionally, we deployed a blocklist to exclude networks that did opt out of previous scans. Scan results will only be published in aggregated form, without any information that could identify individual servers or networks.
@@ 2 Related Work
In 2001, Canetti and Krawczyk [11] established the first model for secure channels, which only requires protection against adversarial insertion of messages. Bellare et al. [3] use stateful authenticated decryption to establish a more robust model, which, as the authors state, should also prevent replay and out-of-order delivery attacks. They use implicit sequence numbers as a shared state between sender and receiver. Paterson et al. [35] define stateful length-hiding authenticated encryption (sLHAE) to model the TLS record layer as a secure channel. This definition was used in [21, 22] to define authenticated and confidential channel establishment (ACCE) to analyze the combination of the TLS handshake and record layer. These steps to formalize secure channels were summarized by Fischlin et al. in [15]. Our attacks show that SSH BPP, when instantiated with ChaCha20-Poly1305, CBC-EtM, or CTR-EtM does not provide integrity of plaintext streams (INT-PST) as defined there.
Formally, SSH BPP security was modelled as stateful decryption [3, 36, 1]. Implicitly, this state was associated with SSH sequence numbers, and it was assumed that this state cannot be manipulated by an adversary. These models can be extended in two directions: (1) Include a broader definition of state. By including chained IVs, key stream state and GCM invocation counters, these models can be used to show why certain cipher modes resist our attacks, and that they indeed achieve INT-PST security. (2) Introduce a novel adversarial query ModifyState to model the attacks described here.
@@
The most severe attack on SSH was presented by Albrecht, Paterson, and Watson [2] in 2009. It exploited the encrypted length field, using the length of the ciphertext accepted by the server from the network as a decryption oracle for a ciphertext block. In [36], this peculiarity of the BPP was formalized, and in [1] a variant of this attack was presented. Other attacks on SSH include a timing attack on SSH keystrokes by Song, Wagner, and Tian [40], a theoretical attack on SSH CBC cipher modes by Wei Dai [12], and a SHA-1 chosen prefix collision attack on the handshake transcript by Bhargavan and Leurent [7]. The weakness of some SSH host keys presented by Heninger et al. [18] was caused by a lack of entropy and faulty implementations and is not an inherent weakness of the protocol.
@@
The SSH handshake was analyzed in [42, 6]. [3] presents a generic security model for SSH BPP, and [36] a specific, more detailed one for SSH-CTR. Albrecht et al. [1] analyze encryption modes in SSH, provide deployment statistics for them, and present a variant of the attack on the length field from [2]. They also include security statements for using ChaCha20-Poly1305, generic Encrypt-then-MAC, and AES-CTR in SSH, claiming the indistinguishability and integrity of the ciphertext. Careful analysis of their proofs reveals some assumptions that do not hold. In particular, they assume that the sequence counter is initialized to 0, which is false for ChaCha20-Poly1305 in SSH. They also assume that the symmetric encryption is a pseudorandom permutation, which is not valid for CBC with IV-chaining. This latter assumption is not apparent from the paper, which omits the proof, but it is discussed carefully in [17] by Hansen, one of the authors.

### v2

# Terrapin Attack: Breaking SSH Channel Integrity By Sequence Number Manipulation

@@ Abstract
The SSH protocol provides secure access to network services, particularly remote terminal login and file transfer within organizational networks and to over 15 million servers on the open internet. SSH uses an authenticated key exchange to establish a secure channel between a client and a server, which protects the confidentiality and integrity of messages sent in either direction. The secure channel prevents message manipulation, replay, insertion, deletion, and reordering. At the network level, SSH uses the Binary Packet Protocol over TCP.
In this paper, we show that as new encryption algorithms and mitigations were added to SSH, the SSH Binary Packet Protocol is no longer a secure channel: SSH channel integrity (INT-PST, aINT-PTXT, and INT-sfCTF) is broken for three widely used encryption modes. This allows prefix truncation attacks where encrypted packets at the beginning of the SSH channel can be deleted without the client or server noticing it. We demonstrate several real-world applications of this attack. We show that we can fully break SSH extension negotiation (RFC 8308), such that an attacker can downgrade the public key algorithms for user authentication or turn off a new countermeasure against keystroke timing attacks introduced in OpenSSH 9.5. Further, we identify an implementation flaw in AsyncSSH that, together with prefix truncation, allows an attacker to redirect the victim’s login into a shell controlled by the attacker.
We also performed an internet-wide scan for affected encryption modes and support for extension negotiation. We find that 71.6% of SSH servers support a vulnerable encryption mode, while 63.2% even list it as their preferred choice.
We identify two root causes that enable these attacks: First, the SSH handshake supports optional messages that are not authenticated. Second, SSH does not reset message sequence numbers when activating encryption keys. Based on this analysis, we propose effective and backward-compatible changes to SSH that mitigate our attacks.
@@ 1 Introduction
While TLS is commonly used to secure user-facing protocols such as web, email, or FTP, SSH is used by administrators to deploy and maintain these servers, often with high privilege (root) access and a large attack surface for lateral movement within an organization’s infrastructure. SSH was developed by Tatu Ylonen in 1995 as a secure alternative to telnet and rlogin/rcp and has since become a critical component of internet security.
@@
An SSH connection between a client and a server begins with the Transport Layer Protocol [33], which defines the handshake messages for key exchange and server authentication and how messages are exchanged over TCP using the Binary Packet Protocol (BPP). After the handshake, SSH provides a secure channel [^1] We note that by channel, we refer to the integrity-protected, encrypted byte stream at the transport level, not the SSH data channels from the connection protocol. Multiple SSH data channels can be multiplexed over the same secure transport channel. for application data. At the application level, the client chooses a sequence of services to run. In practice, the client will run precisely two services: the Authentication Protocol [30] for user authentication with a password or public key, followed by the Connection Protocol [31] for the bulk of SSH’s features like terminal sessions, port forwarding, and file transfer.
@@ 1.1 SSH Channel Security
Figure 1: Typical SSH handshake using a finite-field Diffie-Hellman key exchange. Included sequence numbers are implicit and maintained by the BPP. $\mathsf{Snd}$ denotes the counter for sent packets and $\mathsf{Rcv}$ for received packets. Sequence numbers verified using authenticated encryption are in bold.
In this work, we focus on the integrity of the SSH handshake and the resulting secure channel, as shown in Figure 1. After an initial exchange of version information directly over TCP, the BPP exchanges packets, each containing precisely one message. Initially, the BPP is used without encryption or authentication for the duration of the handshake until the NewKeys message. Afterward, the encryption and authentication keys are used to form a secure channel, intending to protect the confidentiality and integrity of the ordered stream of all following messages. Note that technically, the secure channel consists of two separate cipher streams, one for each direction, and that the order of message arrival is only guaranteed for each direction separately.
@@
As SSH is an interactive protocol, the integrity of each packet must be verified when it is received so that it can be promptly processed. For this, the BPP appends a Message Authentication Code (MAC) to each packet. A cipher mode and a MAC form an authenticated encryption scheme [5]. SSH historically uses Encrypt-and-MAC (EaM), where the MAC is computed over the plaintext, but this is vulnerable to oracle attacks [2]. Later, Encrypt-then-MAC (EtM) was added, where the MAC is computed over the ciphertext instead. SSH has recently adopted the AEAD ciphers AES-GCM and ChaCha20-Poly1305, where ciphertext integrity is built into the encryption scheme [43].
@@ A Trivial Example: Suffix Truncation Attacks
Note that a per-packet MAC cannot fully protect the channel’s integrity, as packets are verified and decrypted before the end of communication has been seen. This allows for a trivial suffix truncation attack, where the attacker interrupts the message flow at some point during the communication. This is an inherent limitation of interactive protocols and an accepted trade-off in the design of SSH, but also, e.g., the TLS Record Layer. Although this attack cannot be prevented, it can be detected by requiring “end-of-communication” messages as the last messages in both directions. TLS defines a “close_notify” alert for this purpose [42]. Although SSH also defines a Disconnect message to indicate the end of the secure channel, this message is optional, unidirectional, and not described as security-critical in the standard.
@@
If the MAC was only computed over the payload of each packet, an attacker could still delete, replay, or reorder packets. Therefore, a sequence number is included in the MAC computation, corresponding to the position of the message in the stream. Each peer maintains two counters (starting at 0), one for each direction. The $\mathsf{Snd}$ counter is incremented after a packet has been sent, and the $\mathsf{Rcv}$ counter is incremented after a received packet has been processed. Once the secure channel has been established, the current value of $\mathsf{Snd}$ is used to compute the MAC of an outgoing packet, and the current value of $\mathsf{Rcv}$ is used to verify the MAC of an incoming packet. If packets in the secure channel are deleted, replayed, or reordered, the sequence numbers get out of sync, and MAC verification will fail.
@@
For TLS, the security guarantees of the Record Layer were formalized as stateful length-hiding encryption [39], with the state mainly consisting of the implicit sequence number. The security of the BPP and implicit sequence numbers was analyzed by Bellare et al. in [4] and later refined and extended by Paterson and Watson [40] and Albrecht et al. [1]. These works define, in slightly idealized scenarios, the following informal security goal for a secure channel:
When a secure channel between A and B is used, the data (or message) stream received by B should be identical to the one sent by A and vice versa (INT-PST, aINT-PTXT in [17]).
Within their idealizations, all three works confirm that the BPP is indeed a secure channel. The difference between the models is that Paterson and Watson [40] also included the encrypted length field of the Encrypt-and-MAC modes, while Albrecht et al. [1] considered the more recent cipher modes ChaCha20-Poly1305, AES-GCM, and generic Encrypt-then-MAC. Our attacks show that the models underlying the proofs in [1] are only partially accurate. We will explain the discrepancies between the proofs and our findings in Section 2.
@@ 1.2 Overview of Our Attacks on SSH
We consider a MitM attacker who can observe, change, delete, or insert bytes at the TCP layer. We do not assume that the attacker can break the confidentiality of the session keys, i.e., the attacker has no information about the derived encryption keys, MAC keys, or IV. However, we do assume that the attacker can determine the length of the messages to be deleted even if the length field is encrypted. We discuss the practicality of this in Section 4.3. The rogue session attack presented in Section 6.2 further assumes the attacker has an account on the same host as the victim.
As for the connection, we assume that the server is correctly authenticated (i.e., the client recognizes the server’s host key) and that a vulnerable encryption mode has been negotiated. See Table 1 for a list of vulnerable encryption modes.
@@
Based on these two key observations, we present a series of novel attacks on SSH that increase in complexity and impact.
@@
We show that an attacker can increase the receive counters of the server and the client by inserting messages into the handshake. Although not required for any of our attacks, we also show that, for some implementations, an attacker can fully control the receive and send counters, setting them to arbitrary values (Section 4.1).
@@ A Prefix Truncation Attack on the BPP
An attacker can use sequence number manipulation to delete a chosen number of packets at the beginning of the secure channel. Neither the client nor the server detects this prefix truncation, consequently breaking the channel integrity of SSH (Section 4.2).
@@
As a practical example, we show an attack that uses prefix truncation to break extension negotiation [9], thereby downgrading the security of the connection. The attacked client might mistakenly believe that the server does not support recent signature algorithms for user authentication or does not implement certain countermeasures to attacks (Section 5.2). Specifically, the attacker can turn off protection against keystroke timing attacks in the recently released OpenSSH 9.5.
@@
As another example, we show two attacks on the AsyncSSH client and server. In the first attack, the victim’s extension info message is replaced with one chosen by the attacker (Section 6.1). For the second attack, the attacker must have a user account on the same server as the victim. The attacker injects a malicious user authentication message so that the victim logs into a shell controlled by the attacker rather than the victim’s shell, thereby giving the attacker complete control over the victim’s terminal input (Section 6.2). These attacks combine prefix truncation with implementation flaws in the AsyncSSH library.
@@
Our attacks critically depend on the SSH encryption mode negotiated between the client and the server. The attack works best with the AEAD cipher ChaCha20-Poly1305 (added in 2013). The attack also works with any EtM mode (added in 2012), although the success probability depends on the cipher mode negotiated. CBC-EtM can be exploited with a significant probability, while the exploitability of CTR-EtM is low. On the other hand, CBC-EaM, CTR-EaM, and GCM modes are not affected. See Section 4.4 for a complete analysis.
In an internet-wide scan, we show that despite these limitations, 71.6% of all SSH servers on the internet support an affected encryption mode, and 63.2% even list it as their preferred choice (Section 7).
@@ 1.3 Our Contributions
  • A first security analysis of SSH extension negotiation, including a novel downgrade attack that disables extension negotiation completely. Thus, support for some public key signature algorithms or, with OpenSSH 9.5, protection against keystroke timing attacks can be disabled.
@@
Proof-of-concept implementations for our attacks and the aggregated results of our internet-wide scan are available under the Apache-2.0 open-source license. See:
@@
We disclosed our findings to 33 vendors of SSH implementations, including OpenSSH and AsyncSSH, in October and November 2023, followed by a public disclosure on December 18th, 2023. As of February 2024, 28 vendors have published patches implementing a backward-compatible countermeasure proposed by OpenSSH. The general protocol flaw has been assigned CVE-2023-48795 (CVSSv3 5.9), while the implementation flaws in AsyncSSH were assigned CVE-2023-46445 (Rogue Extension Negotiation; CVSSv3 5.9) and CVE-2023-46446 (Rogue Session Attack; CVSSv3 6.8). To estimate the adoption rate of the countermeasure, we scanned the IPv4 address space on January 5th, 2024, indicating that more than 3.4M servers were patched.
We provide an opt-out option and an email address for inquiries about our internet-wide scans. Additionally, we employ a block list to exclude networks that opted out of previous scans. Scan results are solely published in aggregated form, without any information that could identify individual servers or networks.
@@ 2 Related Work
In 2001, Canetti and Krawczyk [12] established the first model for secure channels, which only requires protection against adversarial insertion of messages. Paterson et al. [39] defined stateful length-hiding authenticated encryption (sLHAE) to model the TLS record layer as a secure channel. This definition was used in [23, 24] to define authenticated and confidential channel establishment (ACCE) to analyze the TLS handshake and record layer as a whole. Bellare et al. [4] used stateful authenticated decryption to define a security notion for SSH that is directed against replay and out-of-order delivery attacks (INT-sfCTXT). Paterson and Watson [40] later refined this work to cover buffered decryption (INT-BSF-CTXT). Albrecht et al. [1] further refined and extended this definition to cover ciphertext fragmentation attacks more generally (INT-sfCTF). Generalizing the work on TLS and SSH, Fishlin et al. [17] defined, among other notions, plaintext integrity for generic data and (atomic) message streams (INT-PST, aINT-PTXT).
Our attacks show that SSH BPP, when instantiated with ChaCha20-Poly1305, CBC-EtM, or CTR-EtM, does not provide integrity of plaintext or ciphertext (message) streams (INT-PST, aINT-PST, INT-sfCTF) as defined in [1, 17].
@@
The most severe attack on SSH was presented by Albrecht, Paterson, and Watson [2] in 2009. It exploited the encrypted length field, using the length of the ciphertext accepted by the server from the network as a decryption oracle for parts of a ciphertext block. In [40], this peculiarity of the BPP was formalized, and in [1] a variant of this attack was presented. Other attacks on SSH include a timing attack on SSH keystrokes by Song, Wagner, and Tian [45], a theoretical attack on SSH CBC cipher modes by Wei Dai [14], and a SHA-1 chosen prefix collision attack on the handshake transcript by Bhargavan and Leurent [8]. The weakness of some SSH host keys presented by Heninger et al. [20] was caused by a lack of entropy and faulty implementations and is not an inherent weakness of the protocol.
@@
The SSH handshake was analyzed by Williams [47] and Bergsma et al. [7]. Bellare et al. [4] presented a generic security model for SSH BPP, and Paterson and Watson [40] a specific, more detailed one for CTR-EaM. Albrecht et al. [1] included security statements for ChaCha20-Poly1305, generic Encrypt-then-MAC, and AES-CTR in SSH, claiming the indistinguishability and integrity of the ciphertext. Careful analysis of their proofs reveals an essential assumption about SSH sequence numbers that does not hold. In particular, they assume that the sequence counters in the stateful encryption scheme are initialized to 0 on both sides, which is false for the cipher modes affected by our attack. This assumption is not apparent from the paper, which omits the pseudocode for the encryption schemes, but Hansen gives the missing parts in [19] (Alg. ssh-ChaCha20-Poly1305-Gen in Fig. 6.5 and Alg. ssh-fgEtM-Gen in Fig. 6.6 there). We note that this assumption is also present in [4] (Fig. 4 there).
Cadé and Blanchet [11] used the formal verification tool CryptoVerif [10] to prove the security of SSH server authentication and the secrecy of the session key in the computational model. The secrecy of messages in the channel cannot be shown due to the attack in [2]. However, they do mention that due to a limitation in the design of CryptoVerif, it cannot keep mutable internal states such as sequence numbers or counters. In their model, the sequence numbers are passed explicitly as arguments and are, therefore, under the attacker’s control. The authors do not raise the issue of channel integrity. Other computer-aided proofs of server authentication and secrecy of the session key in the symbolic or computational model can be found in [25, 13], which also do not consider the integrity of the secure channel. For an overview of the field of computer-aided cryptography, see [3].